# On the Source Model Key Agreement Problem

### v2

@@ 4 Is the upper bound $I(X;Y)$ tight for the XOR case?
    Show that $S(X;Y\|Z)=S_{\Delta^{*}}(X;Y\|Z)$ for some $\Delta^{*}>0$. Intuitively we expect this to be true because of Proposition 3. We only show this part for the non-interactive setup for the XOR example.
@@
where $P_{X,Y,Z,J,U,V}$ satisfies the following constraint
@@
where in the inner maximum $U$ and $V$ satisfy the cardinality bounds $|\mathcal{V}|\leq|\mathcal{X}||\mathcal{Y}|+1$ and $|\mathcal{U}|\leq|\mathcal{X}||\mathcal{Y}|$.
@@
As observed earlier in computing $\Psi(X;Y\|Z)$ when $Z$ is a function of $(X,Y)$, an optimizer for
@@
is $U=X,Y$ and $V=Z$, which yields the value $I(X,Y;J|Z)$. However, for a positive $\Delta$, this choice of $(U,V)$ may not be feasible in (24).
@@
We now turn to the second task. The following theorem provides a lower bound on $H(Z^{n}|F_{1},F_{2})$ for a capacity achieving code. Some of the manipulations inside the proof are similar to the one given in [18] in the context of a different distributed source coding problem.
@@
Suppose that $X$ and $Y$ are binary random variables, and let $Z=X\oplus Y$. Let
$\Delta_{1}$ $=\min_{U-\!\!\!\!\minuso\!\!\!\!-X-\!\!\!\!\minuso\!\!\!\!-(Y,Z)}H(Z|U)-H(X|U),~{}\Delta_{2}=\min_{V-\!\!\!\!\minuso\!\!\!\!-Y-\!\!\!\!\minuso\!\!\!\!-(X,Z)}H(Z|V)-H(X|V),$ (25)
@@
$\frac{H(Z^{n}|F_{1},F_{2})}{n}\geq\frac{\bar{\Delta}}{2}+\frac{I(K_{A};K_{B}|F_{1},F_{2})}{2n}.$ (27)
@@
$\Delta=\frac{\bar{\Delta}+S_{\text{ni}}(X;Y\|Z)}{2}.$
Then, we can conclude:
$S_{\text{ni}}(X;Y\|Z)=S_{\text{ni},\Delta}(X;Y\|Z).$ (28)
@@
To prove the desired inequality, it suffices to show the following inequalities:
$\frac{1}{n}I(F_{1},F_{2};X^{n},Y^{n})+\frac{2H(Z^{n}|F_{1},F_{2})}{n}$ $\geq H(X,Y)+\min_{U-\!\!\!\!\minuso\!\!\!\!-X-\!\!\!\!\minuso\!\!\!\!-(Y,Z)}H(Z|U)-H(Y|U),$ (29)
$\frac{1}{n}I(F_{1},F_{2};X^{n},Y^{n})+\frac{2H(Z^{n}|F_{1},F_{2})}{n}$ $\geq H(X,Y)+\min_{V-\!\!\!\!\minuso\!\!\!\!-Y-\!\!\!\!\minuso\!\!\!\!-(X,Z)}H(Z|V)-H(X|U),$ (30)
as combining inequalities (29) and (30) gives:
$\frac{1}{n}I(F_{1},F_{2};X^{n},Y^{n})+\frac{2H(Z^{n}|F_{1},F_{2})}{n}\geq H(X,Y)+\bar{\Delta}.$ (31)
We only prove inequality (29), as the proof of (30) is similar. Let $U_{i}=(F_{1},Y^{i-1},Z_{i+1}^{n})$, $V_{i}=(F_{2},X^{i-1},Z_{i+1}^{n})$ be new auxiliary random variables. We have:
$I(F_{1},F_{2};X^{n},Y^{n})+2H(Z^{n}|F_{1},F_{2})$ $=I(F_{1};X^{n},Y^{n})+I(F_{2};Y^{n},X^{n}|F_{1})+2H(Z^{n}|F_{1},F_{2})$
@@
=$\overset{(a)}{=}I(F_{1};X^{n})+H(Y^{n}|F_{1})+H(Z^{n}|F_{1},Y^{n})+I(F_{2};Y^{n}|F_{1},Z^{n})$
$\qquad+H(Z^{n}|F_{1})-H(Y^{n}|F_{1})+I(Y^{n};Z^{n}|F_{1},F_{2})$
=$\overset{(b)}{=}H(X^{n},Y^{n})+I(F_{2};Y^{n}|F_{1},Z^{n})+H(Z^{n}|F_{1})-H(Y^{n}|F_{1})$
$\qquad+I(Z^{n};Y^{n}|F_{1},F_{2})$
@@
=$\overset{(d)}{=}nH(X,Y)+\sum_{i}H(Z_{i}|F_{1},Y^{i-1},Z_{i+1}^{n})-H(Y_{i}|F_{1},Y^{i-1},Z_{i+1}^{n})$
$=nH(X,Y)+\sum_{i}H(Z_{i}|U_{i})-H(Y_{i}|U_{i}),$ (32)
where $(a)$ follows from $H(Z^{n}|F_{1},F_{2},Y^{n})=H(Z^{n}|F_{1},Y^{n})$, which is equivalent with $H(X^{n}|F_{1},F_{2},Y^{n})=H(X^{n}|F_{1},Y^{n})$ which holds due to the Markov chain $F_{2},Y^{n}-\!\!\!\!\minuso\!\!\!\!-X^{n}-\!\!\!\!\minuso\!\!\!\!-F_{1}$, and for $(b)$ we use the following
@@
Also, in the above, $(c)$ follows from the non-negativity of conditional mutual information, and $(d)$ follows from the Marton-Körner identity.
Let $Q$ be the uniform distribution over the set $\{1,2,\cdots,n\}$, independent of $X,Y$ and $F_{1:r}$, and let $X=X_{Q},Y=Y_{Q}$ and $U=(U,Q)$. It is clear that we have the Markov chain $U-\!\!\!\!\minuso\!\!\!\!-X-\!\!\!\!\minuso\!\!\!\!-Y$. Thus, inequality (32) simplifies to the desired bound. ∎
We note that the optimization problems $\min_{U-\!\!\!\!\minuso\!\!\!\!-X-\!\!\!\!\minuso\!\!\!\!-Y}H(Z|U)-H(Y|U)$ and $\min_{V-\!\!\!\!\minuso\!\!\!\!-Y-\!\!\!\!\minuso\!\!\!\!-X}H(Z|V)-H(X|V)$ also arise in many places, including the “modulo 2 sum” problem in [18]. For instance, the following is known:
Example 1 (Proposition $1$ and $2$ of [18]).
Let $P_{X}(0)=x,P_{X|Y}(0|0)=c$ and $P_{X|Y}(1|1)=d$. The value of $\bar{\Delta}$ is 00 if $(1-2c)(1-2d)\leq 0$. Additionally, by setting $x=\frac{\sqrt{d(1-d)}}{\sqrt{d(1-d)}+\sqrt{c(1-c)}}$, the value of $\bar{\Delta}$ is $2H(Z)-H(X,Y)$ if either $c=d$ or $(1-2c)(1-2d)>0$ and $c\neq d$.
5 Acknowgement
Prior to the first author’s involvement in this problem, the second author designated this project as the undergraduate thesis topic for Yicheng Cui. During the initial stages of this research, Cui made some progress on the XOR and AND patterns, employing different (and weaker) techniques than those presented in this work.
The authors would like to also thank Mr. Chin Wa Lau for his help in proving Proposition 2 and for helping Yicheng Cui in his undergraduate thesis.
@@ Appendix A Capacity of a class of source distributions
Assume that $Z=g(X)$ and the channel from $X$ to $Y$ is erasure, i.e., $Y=X$ with probability $1-\epsilon$ and $\{e\}$ otherwise. Note that random variables $X,Y$ and $Z$ do not always form any Markov chain in any order, and the key capacity for this class is not known in the literature.
@@
Consider $V$ constant, and by the functional representation lemma, let $U$ be such that $I(Z;U)=0$ and $H(X|Z,U)=0$. Then, note that $I(U;Z|Y)=0$ since when $Y$ is $X$ it holds and when $Y$ is $e$, it also holds. Then, we have
@@ B.1 Proof of Lemma 1
Note that $I(Z;T|X,Y)=0$ holds trivially when $Z$ is a function of $(X,Y)$. Assume that for given $P_{X,Y}$ there exists a random variable $T$ such that $I(Z;T|X,Y)=I(Z;T)=I(X;Y|T)=0$. Define $Q_{X,Y}^{(t)}=P(X,Y|T=t)$. From $I(X;Y|T)=0$, it follows that $I_{Q^{(t)}}(X;Y)=0$. Additionally, for any $t\in\mathcal{T}$ and $z\in\mathcal{Z}$, we have
$Q^{(t)}(Z=z)=\sum_{x,y\in\mathcal{R}_{z}}Q^{(t)}(X=x,Y=y)=\sum_{x,y\in\mathcal{R}_{z}}P(X=x,Y=y|T=t)=P(Z=z|T=t)\overset{(a)}{=}P(Z=z),$
where $\mathcal{R}_{z}=\left\{(x,y)\in\mathcal{X}\times\mathcal{Y}:z=f(x,y)\right\}$ and $(a)$ follows from $I(Z;T)=0$. Hence, we conclude that $P_{X,Y}\in conv(A)$.
For the converse direction, assume that $P_{X,Y}\in conv(A)$. Therefore, we can represent $P_{X,Y}$ as a convex combination of elements from $A$:
$P_{X,Y}=\sum_{t}\omega_{t}Q_{X,Y}^{(t)},~{}Q_{X,Y}^{(t)}\in A,$ (40)
where $\omega_{t}\geq 0$ and $\sum_{t}\omega_{t}=1$. Define $P(X,Y|T=t)=Q_{X,Y}^{(t)}$. Since $Q^{(t)}\in A$, it follows that $I_{P}(X;Y)=I_{Q^{(t)}}(X;Y)=0$. Next,
$P(Z=z|T=t)=\sum_{x,y\in\mathcal{R}_{z}}P(X=x,Y=y|T=t)=\sum_{x,y\in\mathcal{R}_{z}}Q^{(t)}(X=x,Y=y)=Q(Z=z)\overset{(a)}{=}P(Z=z)$
where $(a)$ holds since $Q^{(t)}\in A$. Thus form above we have $I_{P}(Z;T)=0$. This completes the proof. ∎
@@ B.2 Proof of Theorem 2
The condition $Cov(X,Y)\leq 0$ is equivalent with $p_{00}p_{11}\leq p_{01}p_{10}$.
For the first part, since
$\Psi(X;Y\|Z_{3})\geq\Psi(X;Y\|Z_{1})$
it suffices to show that if $p_{00}p_{11}\leq p_{01}p_{10}$, then $\Psi(X;Y\|Z_{1})=I(X;Y)$ and if $p_{00}p_{11}>p_{01}p_{10}$, then $\Psi(X;Y\|Z_{3})=0.$
Proof of $\Psi(X;Y\|Z_{1})=I(X;Y)$ if $p_{00}p_{11}\leq p_{01}p_{10}$: The condition $p_{00}p_{11}\leq p_{01}p_{10}$ implies that the polynomial
@@
has two non-negative real roots. Let $x_{0}\leq x_{1}$ denote the roots. Note that $x_{0}x_{1}=p_{00}p_{11}$ and $x_{0}+x_{1}=p_{01}+p_{10}$. One can verify that $p_{01}$ belongs to $[x_{0},x_{1}]$. Then, one can find $\lambda\in[0,1]$ such that
@@
Since $\Psi(X;Y\|Z_{1})=\hat{\Psi}(X;Y\|Z_{1})$, to show that $\hat{\Psi}(X;Y\|Z_{1})=I(X;Y)$, we use Theorem 1. It suffices to identify a $T$ such that $I(T;Z_{1}|X,Y)=I(T;Z_{1})=I(X;Y|T)=0$. The condition $I(T;Z_{1}|X,Y)=0$ holds for any $P_{T|X,Y}$ because $Z_{1}$ is a function of $(X,Y)$. Let $T$ be a binary random variable satisfying $P(T=1)=\lambda$ and $P(T=0)=1-\lambda$ and
$P(X=1,Y=1|T=t)=p_{11},\qquad t\in\{0,1\}$
$P(X=0,Y=0|T=t)=p_{00},\qquad t\in\{0,1\}$
$P(X=0,Y=1|T=0)=P(X=1,Y=0|T=1)=x_{1}.$
$P(X=0,Y=1|T=1)=P(X=1,Y=0|T=0)=x_{0}$
@@
Proof of $\Psi(X;Y\|Z_{3})=0$ if $p_{00}p_{11}>p_{01}p_{10}$: we identify a binary random variable $J^{*}$ such that $I(X;Y|J^{*})=I(X,Y;J^{*}|Z_{3})=0$. This would imply that
@@
$\alpha=\dfrac{p_{00}}{p_{00}-(p_{00}p_{11}-p_{01}p_{10})}.$
Under condition $p_{00}p_{11}-p_{01}p_{10}>0$ we have $\alpha\geq 1$. Suppose that
$P_{X,Y|J^{*}=0}=\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$
and
$P_{X,Y|J^{*}=1}=\begin{bmatrix}\alpha p_{00}&\alpha p_{01}\\ \alpha p_{10}&1-\alpha(1-p_{11})\end{bmatrix}$
@@
For the second part, we consider two cases: if $p_{00}p_{11}\leq p_{01}p_{10}$, we use the fact that $\Psi(X;Y\|Z_{1})=I(X;Y)$ and
@@
$\hat{Z}_{1}=\begin{cases}0&(X,Y)=(0,1)\\ 1&(X,Y)=(1,0)\\ 2&(X,Y)\in\{(0,0),(1,1)\}\end{cases}$
Then, by symmetry, from the first part we deduce that in the case of $p_{00}p_{11}\leq p_{01}p_{10}$, we have $\Psi(X;Y\|\hat{Z}_{1})=I(X;Y)$. Since
$\Psi(X;Y\|Z_{2})\geq\Psi(X;Y\|\hat{Z}_{1})=I(X;Y)$
we get the desired result. ∎

### v3

@@ 4 Is the upper bound $I(X;Y)$ tight for the XOR case?
    Show that $S(X;Y\|Z)=S_{\Delta^{*}}(X;Y\|Z)$ for some $\Delta^{*}>0$. Intuitively, we expect this to be true because of Proposition 3. We only show this part for the non-interactive setup for the XOR example.
@@
where $P_{X,Y,Z,J,U,V}$ satisfies the following constraint:
@@
and in the inner maximum, $U$ and $V$ satisfy the cardinality bounds $|\mathcal{V}|\leq|\mathcal{X}||\mathcal{Y}|+1$ and $|\mathcal{U}|\leq|\mathcal{X}||\mathcal{Y}|$.
@@
As observed earlier, in computing $\Psi(X;Y\|Z)$ when $Z$ is a function of $(X,Y)$, an optimizer for
@@
is $U=(X,Y)$ and $V=Z$, which yields the value $I(X,Y;J|Z)$. However, for a positive $\Delta$, this choice of $(U,V)$ may not be feasible in (30). The constraint
$I(V;Z)\leq I(V;J)+H(Z)-\Delta$
for the choice of $V=Z$ reduces to
$\Delta\leq I(Z;J).$
In particular, when $J$ is nearly a constant random variable, $I(Z;J)\approx 0$, which implies that $\Delta$ must be close to zero for $V=Z$ to be admissible. The second task is to prevent $\Delta$ from being close to zero.
@@
We now turn to the second task. The following theorem provides a lower bound on $H(Z^{n}|F_{1},F_{2})$ for a capacity-achieving code. Some of the manipulations inside the proof are similar to those given in [27] in the context of a different distributed source coding problem (the modulo-two sum problem of Körner and Marton).
@@
Suppose that $X$ and $Y$ are binary random variables, and let $Z=X\oplus Y$. Define
$\Delta_{1}$ $=\min_{U-\!\!\!\!\minuso\!\!\!\!-X-\!\!\!\!\minuso\!\!\!\!-(Y,Z)}\bigl{[}H(Z|U)-H(X|U)\bigr{]},~~\Delta_{2}=\min_{V-\!\!\!\!\minuso\!\!\!\!-Y-\!\!\!\!\minuso\!\!\!\!-(X,Z)}\bigl{[}H(Z|V)-H(X|V)\bigr{]},$ (31)
@@
$\frac{1}{n}H(Z^{n}|F_{1},F_{2})\geq\frac{\bar{\Delta}}{2}+\frac{I(K_{A};K_{B}|F_{1},F_{2})}{2n}.$ (33)
@@
$\Delta=\frac{\bar{\Delta}+S_{\mathrm{ni}}(X;Y\|Z)}{2}.$
Then,
$S_{\mathrm{ni}}(X;Y\|Z)=S_{\mathrm{ni},\Delta}(X;Y\|Z).$ (34)
@@
To prove the desired inequality, it suffices to show the following bounds:
$\frac{1}{n}I(F_{1},F_{2};X^{n},Y^{n})+\frac{2}{n}H(Z^{n}|F_{1},F_{2})$ $\geq H(X,Y)+\min_{U-\!\!\!\!\minuso\!\!\!\!-X-\!\!\!\!\minuso\!\!\!\!-(Y,Z)}[H(Z|U)-H(Y|U)],$ (35)
$\frac{1}{n}I(F_{1},F_{2};X^{n},Y^{n})+\frac{2}{n}H(Z^{n}|F_{1},F_{2})$ $\geq H(X,Y)+\min_{V-\!\!\!\!\minuso\!\!\!\!-Y-\!\!\!\!\minuso\!\!\!\!-(X,Z)}[H(Z|V)-H(X|V)].$ (36)
Combining (35) and (36) yields
$\frac{1}{n}I(F_{1},F_{2};X^{n},Y^{n})+\frac{2}{n}H(Z^{n}|F_{1},F_{2})\geq H(X,Y)+\bar{\Delta}.$ (37)
We only prove inequality (35) since (36) is analogous.
Let $U_{i}=(F_{1},Y^{i-1},Z_{i+1}^{n}),\quad V_{i}=(F_{2},X^{i-1},Z_{i+1}^{n})$ be new auxiliary random variables. We have:
$I(F_{1},F_{2};X^{n},Y^{n})+2H(Z^{n}|F_{1},F_{2})$
$=I(F_{1};X^{n},Y^{n})+I(F_{2};Y^{n},X^{n}|F_{1})+2H(Z^{n}|F_{1},F_{2})$
@@
$\overset{(a)}{=}I(F_{1};X^{n})+H(Y^{n}|F_{1})+H(Z^{n}|F_{1},Y^{n})+I(F_{2};Y^{n}|F_{1},Z^{n})$
$\quad+H(Z^{n}|F_{1})-H(Y^{n}|F_{1})+I(Y^{n};Z^{n}|F_{1},F_{2})$
$\overset{(b)}{=}H(X^{n},Y^{n})+I(F_{2};Y^{n}|F_{1},Z^{n})+H(Z^{n}|F_{1})-H(Y^{n}|F_{1})+I(Z^{n};Y^{n}|F_{1},F_{2})$
@@
$\overset{(d)}{=}nH(X,Y)+\sum_{i}H(Z_{i}|F_{1},Y^{i-1},Z_{i+1}^{n})-H(Y_{i}|F_{1},Y^{i-1},Z_{i+1}^{n})$
$=nH(X,Y)+\sum_{i}\left[H(Z_{i}|U_{i})-H(Y_{i}|U_{i})\right],$ (38)
where $(a)$ follows from $H(Z^{n}|F_{1},F_{2},Y^{n})=H(Z^{n}|F_{1},Y^{n})$, which is equivalent to $H(X^{n}|F_{1},F_{2},Y^{n})=H(X^{n}|F_{1},Y^{n})$ which holds due to the Markov chain chain $F_{2},Y^{n}-\!\!\!\!\minuso\!\!\!\!-X^{n}-\!\!\!\!\minuso\!\!\!\!-F_{1}$, and for $(b)$ we use the following
@@
Also, in the above, $(c)$ follows from the non-negativity of conditional mutual information and $(d)$ follows from the Marton-Körner identity.
Let $Q$ be a random variable uniformly distributed over the set $\{1,2,\ldots,n\}$, and let it be independent of the variables $X$, $Y$, and $F_{1:r}$. Let $X=X_{Q},\quad Y=Y_{Q}$ and $U=(U_{Q},Q)$. It is clear that we have the Markov chain $U-\!\!\!\!\minuso\!\!\!\!-X-\!\!\!\!\minuso\!\!\!\!-Y$. Thus, (38) simplifies to the desired bound. ∎
Remark 10.
We note that the optimization problems $\min_{U-\!\!\!\!\minuso\!\!\!\!-X-\!\!\!\!\minuso\!\!\!\!-Y}H(Z|U)-H(Y|U)\quad\text{and}\quad\min_{V-\!\!\!\!\minuso\!\!\!\!-Y-\!\!\!\!\minuso\!\!\!\!-X}H(Z|V)-H(X|V)$ also arise in other contexts, including the ”modulo 2 sum” problem studied in [27]. For instance, the following is known:
Example 1 (Propositions 1 and 2 of [27]).
Let $P_{X}(0)=x$, $P_{X|Y}(0|0)=c$, and $P_{X|Y}(1|1)=d$. Then, $\bar{\Delta}=0$ if $(1-2c)(1-2d)\leq 0$. Moreover, by choosing
$x=\frac{\sqrt{d(1-d)}}{\sqrt{d(1-d)}+\sqrt{c(1-c)}},$
the value of $\bar{\Delta}$ is 2H(Z)H(X,Y)2H(Z)-H(X,Y) if either $c=d$ or (12c)(12d)>0(1-2c)(1-2d)>0 and $c\neq d$.
Acknowledgment
Prior to the involvement of the first author in this problem, the second author had designated this project as the undergraduate thesis topic for Yicheng Cui. During the initial stages of the research, Cui made some progress on the XOR and AND patterns, using different (and generally weaker) techniques than those presented in this work.
The authors also thank Mr. Chin Wa Lau for his valuable assistance in proving Proposition 2 and for supporting Yicheng Cui during his undergraduate thesis.
@@ Appendix A Capacity of a Class of Source Distributions
Assume that $Z=g(X)$ and the channel from $X$ to $Y$ is an erasure channel, i.e., $Y=X$ with probability $1-\epsilon$ and $\{e\}$ otherwise. Note that the random variables $X$, $Y$, and $Z$ do not necessarily form a Markov chain in any order, and the key capacity for this class is not known in the literature.
@@
Consider $V$ as constant, and by the functional representation lemma, let $U$ be such that $I(Z;U)=0$ and $H(X|Z,U)=0$. Then, note that $I(U;Z|Y)=0$ since, when $Y$ is $X$, it holds, and when $Y$ is $e$, it also holds. Then, we have
@@
The significance of this example lies in demonstrating that $S(X;Y\|Z)$ equals the one-way communication rate, despite $X,Y,Z$ not forming a Markov chain in any arrangement.
@@ B.1 Proof of Lemma 1
Note that $I(Z;T|X,Y)=0$ holds trivially when $Z$ is a function of $(X,Y)$. Assume that, for a given $P_{X,Y}$, there exists a random variable $T$ such that $I(Z;T|X,Y)=I(Z;T)=I(X;Y|T)=0$. Define $Q_{X,Y}^{(t)}=P(X,Y|T=t)$. From $I(X;Y|T)=0$, it follows that $I_{Q^{(t)}}(X;Y)=0$. Additionally, for any $t\in\mathcal{T}$ and $z\in\mathcal{Z}$, we have
$Q^{(t)}(Z=z)$ $=\sum_{x,y\in\mathcal{R}_{z}}Q^{(t)}(X=x,Y=y)$
$=\sum_{x,y\in\mathcal{R}_{z}}P(X=x,Y=y|T=t)$
$=P(Z=z|T=t)\overset{(a)}{=}P(Z=z),$
where $\mathcal{R}_{z}=\left\{(x,y)\in\mathcal{X}\times\mathcal{Y}:z=f(x,y)\right\}$, and $(a)$ follows from $I(Z;T)=0$. Hence, we conclude that $P_{X,Y}\in\mathrm{conv}(\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\mathcal{A}})$.
[…]
$P_{X,Y}=\sum_{t}\omega_{t}Q_{X,Y}^{(t)},\quad Q_{X,Y}^{(t)}\in\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\mathcal{A}},$ (46)
where $\omega_{t}\geq 0$ and $\sum_{t}\omega_{t}=1$. Define $P(X,Y|T=t)=Q_{X,Y}^{(t)}$. Since $Q^{(t)}\in\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\mathcal{A}}$, it follows that $I_{P}(X;Y)=I_{Q^{(t)}}(X;Y)=0$. Next,
$P(Z=z|T=t)$ $=\sum_{x,y\in\mathcal{R}_{z}}P(X=x,Y=y|T=t)$
$=\sum_{x,y\in\mathcal{R}_{z}}Q^{(t)}(X=x,Y=y)$
$=Q(Z=z)\overset{(a)}{=}P(Z=z),$
where $(a)$ holds since $Q^{(t)}\in\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\mathcal{A}}$. Thus, from above, we have $I_{P}(Z;T)=0$. This completes the proof. ∎
@@ B.2 Proof of Theorem 2
The condition $\mathrm{Cov}(X,Y)\leq 0$ is equivalent to $p_{00}p_{11}\leq p_{01}p_{10}$.
Consider the first part: we claim that
$\Psi(X;Y\|Z_{3})\geq\Psi(X;Y\|Z_{1}).$ (47)
When $Z$ is a function of $(X,Y)$ we can write
$\Psi(X;Y\|Z)$ $=\min_{J}I(X;Y|J)+I(X,Y;J|Z)$
$=\min_{J}I(X;Y|J)+I(X,Y;J)-I(J;Z).$
Since $Z_{3}$ is a function of $Z_{1}$, we get (47). Since
$\Psi(X;Y\|Z_{3})\geq\Psi(X;Y\|Z_{1}),$
it suffices to show that, if $p_{00}p_{11}\leq p_{01}p_{10}$, then $\Psi(X;Y\|Z_{1})=I(X;Y)$, and if $p_{00}p_{11}>p_{01}p_{10}$, then $\Psi(X;Y\|Z_{3})=0$.
Proof of $\Psi(X;Y\|Z_{1})=I(X;Y)$ when $p_{00}p_{11}\leq p_{01}p_{10}$: The condition $p_{00}p_{11}\leq p_{01}p_{10}$ implies that the polynomial
@@
has two non-negative real roots. Let $x_{0}\leq x_{1}$ denote these roots. Note that $x_{0}x_{1}=p_{00}p_{11}$ and $x_{0}+x_{1}=p_{01}+p_{10}$. One can verify that $p_{01}\in[x_{0},x_{1}]$. Then, there exists $\lambda\in[0,1]$ such that
@@
Since $\Psi(X;Y\|Z_{1})=\hat{\Psi}(X;Y\|Z_{1})$, to show that $\hat{\Psi}(X;Y\|Z_{1})=I(X;Y)$, we use Theorem 1. It suffices to identify a $T$ such that $I(T;Z_{1}|X,Y)=I(T;Z_{1})=I(X;Y|T)=0$. The condition $I(T;Z_{1}|X,Y)=0$ holds for any $P_{T|X,Y}$ because $Z_{1}$ is a function of $(X,Y)$.
Let $T$ be a binary random variable satisfying $P(T=1)=\lambda$ and $P(T=0)=1-\lambda$, and
$P(X=1,Y=1|T=t)=p_{11},\quad t\in\{0,1\},$
$P(X=0,Y=0|T=t)=p_{00},\quad t\in\{0,1\},$
$P(X=0,Y=1|T=0)=P(X=1,Y=0|T=1)=x_{1},$
$P(X=0,Y=1|T=1)=P(X=1,Y=0|T=0)=x_{0}.$
@@
Proof of $\Psi(X;Y\|Z_{3})=0$ when $p_{00}p_{11}>p_{01}p_{10}$: We identify a binary random variable $J^{*}$ such that $I(X;Y|J^{*})=I(X,Y;J^{*}|Z_{3})=0$. This would imply that
@@
$\alpha=\frac{p_{00}}{p_{00}-(p_{00}p_{11}-p_{01}p_{10})}.$
Under the condition $p_{00}p_{11}-p_{01}p_{10}>0$, we have $\alpha\geq 1$. Suppose that
$P_{X,Y|J^{*}=0}=\begin{bmatrix}0&0\\ 0&1\end{bmatrix},\quad P_{X,Y|J^{*}=1}=\begin{bmatrix}\alpha p_{00}&\alpha p_{01}\\ \alpha p_{10}&1-\alpha(1-p_{11})\end{bmatrix}$,
@@
Second part: We consider two cases. If $p_{00}p_{11}\leq p_{01}p_{10}$, we use the fact that $\Psi(X;Y\|Z_{1})=I(X;Y)$ and
@@
$\hat{Z}_{1}=\begin{cases}0&(X,Y)=(0,1),\\ 1&(X,Y)=(1,0),\\ 2&(X,Y)\in\{(0,0),(1,1)\}.\end{cases}$
By symmetry, from the first part we deduce that, in the case $p_{00}p_{11}\leq p_{01}p_{10}$, we have $\Psi(X;Y\|\hat{Z}_{1})=I(X;Y)$. Since
$\Psi(X;Y\|Z_{2})\geq\Psi(X;Y\|\hat{Z}_{1})=I(X;Y),$
we obtain the desired result. ∎